{ProBlow #4} For this Given Gneust Fan ilt). 60 k h 30h 8 I to
Answer #1

initial conditions : for tto C cap is open ) : -vc61+2=0 VC 6 )= PV 60k * м Кл i Yi2 + 60kn Vco) 1.Uc6 )=2V1 SV L (OR) (OR moJn s-Domain : GOKN v iz 30kn white V GORN jy is nodal at node vi itiz+ i3 =0 is = 165) | is) | 1 t s V- Slov V-215 con 60K 30240m 6+4 + 120ms 5 rooms +1+ Goms] 360ms +4 5 ( 180m 5 +1) Erein 360m 5+ 4 GOK is) - - using - Bom StToomstu partial fractionA = 4 = 66.67 el B= (1.90ms +1 ) [s? | s2-1/100m = (180ms A11 (360ms +12 GOK $ (120996+1) |-- - 360m cicom) + sok t lizom) 3= 15! r£9.99 alu 120msti ; G6. 671 16.67 1 a st 4120mce (t) a 1 I e un 16 는 sta inverse Trang form Takingist) = 66.67 .414) + - 16.67 u ē 1200 U CH) (66. 670 + 16.67 - tl 120m Ju(t) ē ar UA.